# MonarchAttention: Zero-Shot Conversion to Fast, Hardware-Aware Structured Attention

### v1

@@ 4 Experiments
Finally, we validate that the computational/IO complexity reduction achieved by MonarchAttention translates into actual speed-ups on the NVIDIA A40, a modern GPU. We implement the pseudo-code described in Section 3 as four separate Triton kernels and compare it against the fastest available implementations of FlashAttention-2 – either the Triton implementation or PyTorch’s scaled_dot_product_attention, which calls the CUDA backend for FlashAttention-2. Using a fixed batch size $E$, number of heads $H$, and head dimension $d$, we sweep the input sequence length $N$ and compare the run-times of FlashAttention-2 and MonarchAttention (with $b=\sqrt{N}$ and $T=1$) in Figure 6 (left). As sequence length increases, MonarchAttention consistently outperforms FlashAttention-2, notably achieving up to 8.2×8.2$\times$8.2 × speed-up with $N=16384$. To highlight gains for shorter sequences, we implement MonarchAttention as a single fully-fused Triton kernel, with a single thread block computing a single head. For fixed sequence length $N=256$, number of heads $H$, and head dimension $d$, we sweep the batch size $E$ and compare the run-time of the fully-fused MonarchAttention kernel against FlashAttention-2 in Figure 6 (right). With smaller batch sizes, we have low utilization of hardware, since we compute a single head with a single thread block. However, as we increase the batch size, MonarchAttention achieves up to 1.4×1.4$\times$1.4 × speed-up over FlashAttention-2.
Figure 6: Run-times of MonarchAttention and FlashAttention-2 across various sequence lengths. Normalized runtime ($1=\mbox{slowest},0=\mbox{fastest}$) of MonarchAttention and FlashAttention-2 on NVIDIA A40 GPU. Left: sweep of sequence length $N$ with $E=1$, $H=12$, and $d=64$. Right: sweep of batch size $E$ with $N=256$, $H=12$, and $d=64$.
@@ 5 Conclusion
To conclude, we discuss several limitations and future directions for this work. First, the implementation of MonarchAttention can be improved, particularly with more recent GPUs having expanded capabilities such as distributed shared memory found on Hopper architectures. Next, while we have presented MonarchAttention as a direct replacement for softmax attention with no additional training, in theory it can also accelerate training from scratch or achieve better results by converting existing models with some fine-tuning. Moreover, MonarchAttention currently does not support causal masking, which could accelerate training of language models trained with next-token prediction. Finally, we believe that viewing fundamental operations such as softmax through their variational form is a powerful idea that can be generalized, allowing for more generic structured approximations beyond Monarch approximations to softmax attention.
@@ Appendix B Monarch Background
We provide an example of the transpose permutation $\bm{P}$ in Section 2. Recall that applying $\bm{P}\in\mathbb{R}^{N\times N}$ to a vector $\bm{x}\in\mathbb{R}^{N}$ corresponds to row-major reshaping $\bm{x}$ to $\mathbb{R}^{m\times b}$, transposing to $\mathbb{R}^{b\times m}$, then row-major flattening back to $\mathbb{R}^{N}$. This is equivalent to applying a permutation matrix whose $(i+1)$th row is given by $\bm{e}_{\sigma(i)+1}$ where
@@ C.1 Updates
$\bm{\beta}_{R,kji}$ $=\sum_{v}\overline{\bm{K}}_{kiv}\sum_{l}(\bm{L}_{jkl}\overline{\bm{Q}}_{jlv}),\quad\bm{\gamma}_{R,kj}=\sum_{l}\bm{L}_{jkl}\log(\bm{L}_{jkl}),\quad\bm{\gamma}_{R,kj}=\sum_{l}\bm{L}_{jkl}.$
@@
$\log\bm{R}_{kji}=\frac{\bm{\beta}_{R,kji}-\bm{\gamma}_{R,kj}-\bm{\lambda}_{R,kj}}{\bm{c}_{R,kj}}-1\implies\bm{R}_{kji}=\frac{\exp(\bm{\beta}_{R,kji}/\bm{c}_{R,kj})}{\exp(\bm{\gamma}_{R,kj}/\bm{c}_{R,kj}+\bm{\lambda}_{kj}^{R}/\bm{c}_{R,kj}+1)}.$
Finally, from the constraint $\sum_{i}\bm{R}_{kji}=1$, we must have $\exp(\bm{\gamma}_{R,kj}/\bm{c}_{R,kj}+\bm{\lambda}_{kj}^{R}/\bm{c}_{R,kj}+1)=\sum_{i}\exp(\overline{\bm{\alpha}}_{kji}^{R}/\bm{\gamma}_{kj}^{R})$, which gives the final closed form update:
@@ C.2 Naïve Algorithm
We provide pseudo-code for the naive version of MonarchAttention in Figure 7. This is a direct implementation of alternating maximization for finding the Monarch factors, ignoring concerns about memory and IO. We alternate between updating $\bm{L}$ and $\bm{R}$ as described above for $T$ iterations. Specifically, we highlight the choices to (1) initialize $\bm{L}$ to be block identity, and (2) update $\bm{R}$ before $\bm{L}$ in each iteration.
{minted}
[frame=single,linenos,fontsize=]python # Q: array of size (N, d) # K: array of size (N, d) # V: array of size (N, d) # T: number of steps
def monarch_attention(Q, K, V, T): L = stack(b * [eye(m)]) Qb = einshape(”(lj)v-¿jlv”, Q, j=b) Kb = einshape(”(ki)v-¿kiv”, K, i=b)
# Alternating maximization for L, R for t in range(T): # R update aR = einsum(”jkl,jlv-¿kjv”, L, Qb) bR = einsum(”kjv,kiv-¿kji”, aR, Kb) cR = einsum(”jkl-¿kj”, L) R = softmax(bR / cR[:, :, None], axis=2)
# L update aL = einsum(”kji,kiv-¿jkv, R, Kb) bL = einsum(”jkv,jlv-¿jkl”, aL, Qb) cL = einsum(”kji-¿jk”, R * log(R)) L = softmax(bL - cL[:, :, None], axis=1)
# Monarch multiply Vb = einshape(”(ki)v-¿kiv”, V, i=b) Y = einsum(”kji,kiv-¿jkv”, R, Vb) Z = einsum(”jkl,jkv-¿ljv, L, Y) O = einshape(”ljv-¿(lj)v”, Z)
return O
@@ C.3 Padding and Masking
From (5), it is clear that to set all columns of $\bm{M}$ beyond the $N$th column to zero, it is sufficient to set $\bm{R}_{kji}$ to zero whenever $b(k-1)+i>N$. Thus, we simply form the mask $\bm{\omega}\in\mathbb{R}^{m^{\prime}\times b\times b}$ given by
@@ D.1 Baselines
    $\tilde{\bm{F}}=\operatorname{softmax}(\bm{Q}\tilde{\bm{K}}^{\top}),\;\tilde{\bm{B}}=\operatorname{softmax}(\tilde{\bm{Q}}\bm{K}^{\top}),\;\tilde{\bm{A}}=\operatorname{softmax}(\tilde{\bm{Q}}\tilde{\bm{K}}^{\top})$
@@ D.5 Image Generation with DiT
To create the images in Figure 5, we used a random seed of 00 input the same $36$ random Gaussian samples into all three models. To obtain the results in Table 1, we again used a random seed of 00, and generated $50K$ images from each type of model, again using the same $50K$ random samples across all models.

### v2

@@ 4 Experiments
Finally, we validate that the computational/IO complexity reduction achieved by MonarchAttention translates into actual speed-ups on the NVIDIA A40, a modern GPU. We implement the pseudo-code described in Section 3 as four separate Triton kernels and compare it against the fastest available implementations of FlashAttention-2 – either the Triton implementation or PyTorch’s scaled_dot_product_attention, which calls the CUDA backend for FlashAttention-2. Using a fixed batch size $E$, number of heads $H$, and head dimension $d$, we sweep the input sequence length $N$, generate random $\bm{Q},\bm{K},\bm{V}$ matrices, and compare the run-times of FlashAttention-2 and MonarchAttention (with $b=\sqrt{N}$ and $T=1$) in Figure 6 (left). As sequence length increases, MonarchAttention consistently outperforms FlashAttention-2, notably achieving up to $8.2\times$ speed-up with $N=16384$. To highlight gains for shorter sequences, we implement MonarchAttention as a single fully-fused Triton kernel, with a single thread block computing a single head. For fixed sequence length $N=256$, number of heads $H$, and head dimension $d$, we sweep the batch size $E$ and compare the run-time of the fully-fused MonarchAttention kernel against FlashAttention-2 in Figure 6 (right). With smaller batch sizes, we have low utilization of hardware, since we compute a single head with a single thread block. However, as we increase the batch size, MonarchAttention achieves up to $1.4\times$ speed-up over FlashAttention-2.
Figure 6: Run-times of MonarchAttention and FlashAttention-2 across various sequence lengths. Normalized runtime ($1=\mbox{slowest},0=\mbox{fastest}$) of MonarchAttention and FlashAttention-2 on NVIDIA A40 GPU. Left: sweep of sequence length NN with $E=1$, $H=12$, and $d=64$. Right: sweep of batch size $E$ with $N=256$, $H=12$, and $d=64$.
@@ 5 Conclusion
To conclude, we discuss several limitations and future directions for this work.
Autoregressive Transformers.
In this paper, we have primarily focused on zero-shot approximation of attention in encoder-based Transformers, yet most generative language models of interest today are decoder-based autoregressive Transformers. One limitation of MonarchAttention is that it fundamentally cannot be applied to autoregressive generation, since there is no “attention matrix” to approximate – instead, single queries are passed through the attention mechanism in a streaming fashion, and discarded after each decode step. Instead, MonarchAttention could be used during prefill to efficiently compute keys/values for a provided prefix (e.g., long prompt or context). Moreover, MonarchAttention can be applied to emerging non-autoregressive paradigms such as diffusion language models (DLMs) to accelerate language generative models.
Accelerating Training.
While we have presented MonarchAttention as a direct replacement for softmax attention with no additional training, it can also accelerate pre-training or fine-tuning as well. Although MonarchAttention cannot be applied at inference time for autoregressive Transformers, it can be applied to training such models. However it is not obvious how to efficiently incorporate the causal mask into MonarchAttention.
Beyond Monarch & Softmax.
Finally, although MonarchAttention is specifically designed for approximating softmax attention via Monarch, this idea can be generalized to other (possibly dynamic) structured matrix classes, as well as other nonlinear mappings. Monarch, while efficient, allocates the same number of parameters to each uniformly sized block. Yet, in practice, different regions of the attention matrix may require more fine-grained approximation. On the other hand, $\alpha$-entmax mappings (Peters et al., 2019), which generalize softmax, can produce sparser attention matrices, thereby resulting in better approximations via block rank-one type matrices.
@@ Appendix B Monarch Background
We provide an example of the transpose permutation $\bm{P}$ in Section 2. Recall that applying $\bm{P}\in\mathbb{R}^{N\times N}$ to a vector $\bm{x}\in\mathbb{R}^{N}$ corresponds to row-major reshaping $\bm{x}$ to $\mathbb{R}^{m\times b}$, transposing to $\mathbb{R}^{b\times m}$, then row-major flattening back to $\mathbb{R}^{N}$. This is equivalent to applying a permutation matrix whose $(i+1)$-th row is given by $\bm{e}_{\sigma(i)+1}$ where
@@ C.1 Updates
$\bm{\beta}_{R,kji}$ $=\sum_{v}\overline{\bm{K}}_{kiv}\sum_{l}(\bm{L}_{jkl}\overline{\bm{Q}}_{jlv}),\quad\bm{\gamma}_{R,kj}=\sum_{l}\bm{L}_{jkl}\log\bm{L}_{jkl},\quad\bm{c}_{R,kj}=\sum_{l}\bm{L}_{jkl}.$
@@
$\log\bm{R}_{kji}=\frac{\bm{\beta}_{R,kji}-\bm{\gamma}_{R,kj}-\bm{\lambda}_{R,kj}}{\bm{c}_{R,kj}}-1\implies\bm{R}_{kji}=\frac{\exp(\bm{\beta}_{R,kji}/\bm{c}_{R,kj})}{\exp(\bm{\gamma}_{R,kj}/\bm{c}_{R,kj}+\bm{\lambda}_{R,kj}/\bm{c}_{R,kj}+1)}.$
Finally, from the constraint $\sum_{i}\bm{R}_{kji}=1$, we must have $\exp(\bm{\gamma}_{R,kj}/\bm{c}_{R,kj}+\bm{\lambda}_{R,kj}/\bm{c}_{R,kj}+1)=\sum_{i}\exp(\bm{\beta}_{R,kji}/\bm{c}_{R,kj})$, which gives the final closed form update:
@@ C.2 Naïve Algorithm
We provide pseudo-code for the naive version of MonarchAttention in Figure 7. This is a direct implementation of alternating maximization for finding the Monarch factors, where we highlight the choices to (1) initialize $\bm{L}$ to be block identity, and (2) update $\bm{R}$ before $\bm{L}$ in each iteration. We include this here for completeness, so that readers can directly compare this code with the algorithmic steps. The implementation in Figure 4 is a reorganization of Figure 7 where $\bm{L},\bm{R}$ are not materialized in HBM, resulting in FlashAttention-like kernels.
# Q: array of size (N, d)
# K: array of size (N, d)
# V: array of size (N, d)
# T: number of steps
[…]
L = stack(b * [eye(m)])
Qb = einshape(”(lj)v->jlv”, Q, j=b)
Kb = einshape(”(ki)v->kiv”, K, i=b)
[…]
for t in range(T):
# R update
aR = einsum(”jkl,jlv->kjv”, L, Qb)
bR = einsum(”kjv,kiv->kji”, aR, Kb)
cR = einsum(”jkl->kj”, L)
R = softmax(bR / cR[:, :, None], axis=2)
[…]
aL = einsum(”kji,kiv->jkv”, R, Kb)
bL = einsum(”jkv,jlv->jkl”, aL, Qb)
cL = einsum(”kji->jk”, R * log(R))
L = softmax(bL - cL[:, :, None], axis=1)
[…]
Vb = einshape(”(ki)v->kiv”, V, i=b)
Y = einsum(”kji,kiv->jkv”, R, Vb)
Z = einsum(”jkl,jkv->ljv”, L, Y)
O = einshape(”ljv->(lj)v”, Z)
[…]
@@ C.3 Padding and Masking
From (5), it is clear that to set all columns of $\bm{M}$ beyond the $N$-th column to zero, it is sufficient to set $\bm{R}_{kji}$ to zero whenever b(k1)+i>Nb(k-1)+i>N. Thus, we simply form the mask $\bm{\omega}\in\mathbb{R}^{m^{\prime}\times b\times b}$ given by
@@ D.1 Baselines
[…]
@@ D.2 Image Classification with ViT
The choice of $b=14$ for monarch-attention is due to the fact that images are patched in a $14\times 14$ grid. We also apply post-padding as described in Section C.3 for the [CLS] token, since it is appended at the end of the sequence.
@@ D.5 Image Generation with DiT
To create the images in Figure 5, we used a random seed of 0 input the same $36$ random Gaussian samples into all three models. To obtain the results in Table 1, we again used a random seed of 0, and generated $50K$ images from each type of model, again using the same $50K$ random samples across all models.
Appendix E Additional Experimental Results
E.1 Convergence of MonarchAttention
We empirically show that MonarchAttention converges very quickly in the ViT image classification setting described in Section D.2. We measure of the value of the softmax variational objective $f(\bm{M}^{(T)};\bm{Q},\bm{K})$ as defined in (6) in the 0-th layer and $5$-th head for the first few $T$ – the results are shown in Table 3. We see that by $T=2$, the variational objective has already reached a stationary point, with even $T=1$ closing most of the gap. This demonstrates how MonarchAttention is able to reach an accurate solution with even just one step.
Table 3: Convergence of $f(\bm{M}^{(T)};\bm{Q},\bm{K})$.
$T$ 0 11 22 33
$f(\bm{M}^{(T)})$ $418$ $937$ $940$ $940$
E.2 Layer Ablation for RoBERTa Question Answering
We investigate the impact of replacing different layers with MonarchAttention in the RoBERTa question answering setting described in Section D.3. The results are shown in Table 4.
Table 4: F1 score when replacing a given layer of RoBERTa with MonarchAttention.
Layer 0 11 22 33 44 55 66 77 88 99 $10$ $11$
F1 ($\uparrow$) $92.7$ $92.5$ $92.6$ $92.0$ $91.4$ $91.9$ $86.3$ $91.1$ $92.7$ $93.2$ $93.0$ $92.3$
The initial and final layers are more amenable to replacement compared to the middle layers.
E.3 Post-Swap Fine-Tuning for BART Summarization
We demonstrate the improved performance of MonarchAttention when fine-tuning on BART summarization post-replacement. We follow the setting and hyperparameters of Section D.4, but fine-tune with MonarchAttention or Nyströmformer instead of the standard softmax attention for sequence length $N=8192$. The results are shown in Table 5, where we also copy the zero-shot results from the main body for reference.
Table 5: Post-swap fine-tuning vs zero-shot performance on BART summarization.
Type Method Total Attention FLOPs ($10^{9}$) ROUGE-1 ($\uparrow$)
Softmax 618 35.57
Fine-tuned MonarchAttention 31 35.38
Nyströmformer 35 33.94
Zero-shot MonarchAttention 31 34.60
Nyströmformer 35 32.86
We see that MonarchAttention is stable under fine-tuning, and is indeed a strong approximation even beyond the zero-shot setting. In particular, we can almost fully close the gap with MonarchAttention to the original softmax model, while benefiting from more efficient fine-tuning with roughly 5% of the attention FLOPs.
E.4 Layer Ablation for DiT Image Generation
We investigate the impact of replacing different quarters with MonarchAttention in the DiT image generation setting described in Section D.5. The results are shown in Table 6. We see that replacing earlier layers is more favorable than later layers, with the most degradation resulting from replacing the final layers of the DiT.
Table 6: (s)FID scores when replacing different quarters of DiT with MonarchAttention.
Quarter First Second Third Fourth
FID ($\downarrow$) 0.07 0.21 0.34 1.35
sFID ($\downarrow$) 0.14 0.39 0.52 2.23
E.5 Graph Node Classification with GraphGPS
We evaluate the efficacy of MonarchAttention beyond language and image modalities by considering zero-shot conversion of attention in the GPS graph Transformer (GPS-GT) (Rampášek et al., 2022) on the Actor graph dataset (Pei et al., 2020) for node classification. In particular, we convert the attention layer in a single-layer 170K parameter GPS-GT with 2 heads, sequence length $N=7600$, and head dimension $d=48$. We train for 1500 steps with learning rate of $5\times 10^{-4}$ using the Adam optimizer, and also evaluate on the train split [^4] We evaluate on the train split due to low performance on the validation/test sets due to limited data/training from scratch. Since we are primarily interested in the post-train zero-shot approximation capabilities of MonarchAttention, this does not introduce bias in the evaluation.. To evaluate the performance at different FLOP counts, we vary the number of steps $T\in\{1,2,3,4\}$ while fixing $b=128$ for MonarchAttention, and vary the rank $r\in\{128,192,256,320\}$ for Nyströmformer. The results are shown in Figure 8.
Figure 8: Zero-shot conversion of attention layers for graph node classification. We vary hyperparameters to evaluate training accuracy vs. total attention FLOPs across all layers for GPS on the Actor dataset.
[…]